In financial and managerial accounting (14th edition) chapter 17, problem 12E

I'm trying to understand how the overhead costs were calculated. The answer is $285,000 but I'm not sure how that was calculated. Below are the details. Thanks!

Fenwick Corporation's manufacturing and finished goods warehouse facilities burned to the ground on January 31. The loss was fully covered by insurance. The insurance company wanted to know the cost of the inventories destroyed in the fire. The company's accountants gathered the following information:

Direct materials purchased in January - $160,000

Work in process inventory, January 1 - $34,000

Materials Inventory, January 1 - $16,000

Finished Goods Inventory, January 1 - $30,000

Direct Labor Costs in January $190,000

Prime costs charged to jobs in January $294,000

Cost of finished goods available for sale in January - $450,000

Sales revenue earned in January - $500,00

Gross profit as a percentage of January sales- 25%

Manufacturing overhead applied to jobs in January as a % of total conversion costs - 60%

Assume that actual manufacturing overhead was exactly equal to the amount applied to production at the time of the fire.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Fenwick Corporation Amount $
Let Total conversion costs for January 100X
Manufacturing overhead applied to jobs in January as a % of total conversion costs 60%
Manufacturing overhead applied (100X*60%) 60X
So, total conversion costs for January 100X
Less: Manufacturing overhead applied 60X
Direct Labor Costs in January 40X
But it is given that,
Direct Labor Costs in January 190,000.00
So, 40X= 190,000.00
X= 190,000/40=        4,750.00
Manufacturing overhead applied= 60X
Manufacturing overhead applied= 60*4750 285,000.00
Manufacturing overhead is $285,000.